Show the graph and determine if the given function is continuous on each of the given intervals.




1. f(x)=4x²-x+6; (-infinite, 0)



2.f(x)=4/x-5;(-5,5)



3.f(x)=√x-1;(1,+infinite)



4.f(x)=-|-5x|;(-1,+infinite)

1
Expert's answer
2022-03-17T13:38:31-0400

Solution (1)


For the function

"f(x)=4x\u00b2-x+6" over the interva"(-\infty, 0]"


The plot is shown below





From the plot, we can see that "f(x)=4x\u00b2-x+6;" is continouse over the interval "(-\infty, 0]"


Solution (2)


For the function

"f(x)=\frac{4}{x}" over the interval "(-5,5)" the plot is shown below





From the plot we can see that the function "f(x)=\frac{4}{x}" is not continous at "x=0"


Solution (3)


For the function

"f(x)=\sqrt{x-1}" over the interval "(1,\infty)" the plot is shown below





From the plot we can see that the function "f(x)=\sqrt{x-1}" over the interval "(1,\infty)" is continous


Solution (4)


For the function "f(x)=-|-5x|" over the interva"(-1, \infty)"the plot is shown below





From the plot we can see that the function "f(x)=-|-5x|" over the interval "(-1,\infty)" is continous